### v1

@@ 4.2 Identification and Tokenization
25 "taxonomies": [
26 {"system": "UNSPSC", "code": "70131500", "label": "Residential property"}
27 ],
28 "spatialFootprint": {
29 "encoding": "GeoJSON",
30 "geometry": {
31 "type": "Polygon",
32 "coordinates": [[[121.45,31.20],[121.46,31.20],[121.46,31.21],[121.45,31.21],[121.45,31.20]]]
33 },
34 "granularity": "parcel"
35 },
36 "documents": [
37 {"name": "Deed", "hash": "0xabc...", "mediaType": "application/pdf", "issuedBy": "did:web:issuer.org"}
38 ],
39 "relations": [
40 {"relation": "belongsToClass", "target": "did:web:issuer.example.org:class:RE-RESIDENCE"}
41 ],
42 "attributes": [
43 {"name": "floorArea", "value": "120", "unit": "sqm"}
44 ],
45 "custom": {
46 "localPolicyTag": "RWA-POL-01",
47 "iotDeviceIds": ["sensor-xyz-001"]
48 },
@@
51},
52"compliance": {
53 "licenseId": "example-2025-8899",
54 "sellableRegions": [//Jurisdiction Code],
55 "restrictions": ["NoCrossBorderSale"],
56 "effectiveFrom": "2025-01-01",
57 "effectiveTo": "2026-01-01",
58 "regulatorDid": "did:web:regulator.gov",
@@
61},
62"custody": {
63 "custodianDid": "did:web:custody.bank.example",
64 "location": "Vault-example-01",
65 "policy": "ISO-27001+SOC2",
66 "auditCycleDays": 30,
67 "insurancePolicyRef": {"hash": "0xins..."},
@@ 4.3 Cross-Chain Authentication
Consider an RWA asset $a$ that has been tokenized on $C_{1}$, producing a credential $\mathsf{cred}$ anchored by a transaction $tx$ in block $B$. To enable recognition of $a$ on another chain $C_{2}$ without re-executing the full tokenization procedure, we employ SPV as defined in Section 3.3.
@@
Unlike conventional approaches where each blockchain must independently repeat the full verification of decentralized identifiers and verifiable credentials, the SPV-based protocol requires such verification only once on the source chain. Subsequent chains merely validate the compact SPV proof to ensure that the authenticated transaction has indeed been recorded on the source ledger and that the embedded commitment is consistent with the disclosed credential fields. This design eliminates redundant executions of DID resolution, credential signature checks, and revocation list queries across multiple ledgers, thereby reducing both computation and communication overhead while maintaining consistency of authentication results.
@@ 4.4 Efficient Cross-Chain Interactions
In this part, we adopt HTLCs and off-chain channels to construct a cross-chain channel that supports partial settlement of RWAs without requiring channel closure. We first provide the definition of the cross-chain channel and then describe how it can be applied to the settlement of RWAs.
@@
Definition 6 (Cross-Chain Channel)
@@
The above definition shows that settlement of RWAs can be achieved within the channel lifecycle through the $\mathsf{Unlock}$ and $\mathsf{Refund}$ procedures, without requiring channel closure. This feature allows partial settlements to be executed while keeping the channel open, thereby reducing the cost of repeated channel establishment and improving the efficiency of cross-chain interactions.
Cross-Chain Interactions. First, the parties invoke $\Pi.\mathsf{Open}$ by deploying contracts $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$ on $C_{1}$ and $C_{2}$, which establish the channel environment and record deposits. During the lifetime of the channel, they repeatedly apply $\Pi.\mathsf{Update}$ to exchange signed off-chain states describing tentative transfers of different subsets of RWAs. For example, a state $\sigma_{j}$ may specify that $\mathcal{B}$ pays $v_{j}$ units on $C_{1}$ in exchange for batch $S_{j}\subseteq\{a_{1},\dots,a_{m}\}$. These updates remain entirely off-chain, enabling multiple negotiations without incurring blockchain overhead. Once the parties agree on a settlement, $\Pi.\mathsf{Lock}$ is executed: both sides upload the agreed final state to $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$, where hash-locked conditions $h(\rho)$ with timeouts $T_{1}>T_{2}$ are installed to ensure atomicity. If $\mathcal{B}$ reveals the preimage $\rho$ on $\xi^{\mathsf{2}}$ before $T_{2}$, $\Pi.\mathsf{Unlock}$ delivers the agreed union of RWAs $\bigcup_{j}S_{j}$ to $\mathcal{B}$, and allows $\mathcal{S}$ to redeem the corresponding aggregated payment on $\xi^{\mathsf{1}}$ before $T_{1}$. If instead $\mathcal{B}$ does not reveal $\rho$ in time, $\Pi.\mathsf{Refund}$ is triggered: $\xi^{\mathsf{2}}$ returns the RWAs to $\mathcal{S}$ at $T_{2}$, and $\xi^{\mathsf{1}}$ returns the funds to $\mathcal{B}$ after $T_{1}$. Finally, once the session concludes, $\Pi.\mathsf{Close}$ is called to release any remaining deposits and terminate the channel.
@@
In the following section, we demonstrate the feasibility of our scheme by conducting simulation-based experiments.
@@ 5.1 Implementation
Experiments are conducted on a workstation with an Intel Core i9-13980HX CPU and 32 GB RAM, running Windows 11 (24H2). Python 3.12.10 is used to implement and benchmark the VC routines and the SPV algorithm, and the reported execution times are obtained from these Python implementations under multi-threaded settings. Smart-contract experiments are performed on Remix VM (Prague) using Solidity 0.8.30 to measure gas consumption for the HTLC and cross-chain channel workflows. The entire codebase consists of more than 2,000 lines, including the Python implementations for VC/SPV and the Solidity contracts for on-chain evaluation. Our code repository is publicly available at https://github.com/ZhuHaoming2005/xRWA.
@@ 5.2 Performance Evaluation
For the cross-chain channel, handling $n$ cross-chain interactions requires only 917,253 gas in total, since all updates are performed off-chain and only the opening and unlocking operations are executed on-chain. In contrast, the HTLC scheme consumes 465,426 $\times\,n$ gas for $n$ interactions, as each interaction must execute both the lock and unlock operations on-chain. This demonstrates that the cross-chain channel achieves significantly higher efficiency than HTLC, especially when the number of interactions increases.

### v3

@@ 4.3 Identification and Tokenization
25 //Attributes applicable to specific scenarios
26 ...
@@
29},
30"compliance": {
31 "licenseId": "<string>",
32 "sellableRegions": ["<string>"],
33 "restrictions": ["<string>"],
34 "effectiveFrom": "<timestamp>",
35 "effectiveTo": "<timestamp>",
36 "regulatorDid": "<did:example:regulator>",
@@
39},
40"custody": {
41 "custodianDid": "<did:example:custody>",
42 "location": "<string>",
43 "policy": "<string>",
44 "auditCycleDays": <number>,
45 "insurancePolicyRef": "<string>",
@@ 4.4 Cross-Chain Authentication
Consider an RWA $a$ that has been tokenized on $C_{1}$, producing a credential $\mathsf{cred}$ anchored by a transaction $tx$ in block $B$. To enable recognition of $a$ on another chain $C_{2}$ without re-executing the full tokenization procedure, we employ SPV as defined in Section 3.3.
@@
Unlike conventional approaches where each blockchain must independently repeat the full verification of decentralized identifiers and verifiable credentials, the SPV-based protocol requires such verification only once on the source chain. Subsequent chains merely validate the compact SPV proof to ensure that the authenticated transaction has indeed been recorded on the source ledger and that the embedded commitment is consistent with the disclosed credential fields.
@@ 4.5 Efficient Cross-Chain Interactions
In this part, we adopt HTLCs and off-chain channels to construct a cross-chain channel that supports partial settlement of RWAs without requiring channel closures. We first provide the definition of the cross-chain channel and then describe how it can be applied to the settlement of RWAs.
@@
Definition 7 (Cross-Chain Channel).
@@
The above definition shows that settlement of RWAs can be achieved within the channel lifecycle through the $\mathsf{Unlock}$ and $\mathsf{Refund}$ procedures, without requiring channel closures. This feature allows partial settlements to be executed while keeping the channel open, thereby reducing the cost of repeated channel establishment and improving the efficiency of cross-chain interactions.
Cross-Chain Interactions. First, the parties invoke $\Pi.\mathsf{Open}$ by deploying contracts $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$ on $C_{1}$ and $C_{2}$, which establish the channel environment and record deposits. During the lifetime of the channel, they repeatedly apply $\Pi.\mathsf{Update}$ to exchange signed off-chain states describing tentative transfers of different subsets of RWAs. For example, a state $\sigma_{j}$ may specify that $\mathcal{B}$ pays $v_{j}$ units on $C_{1}$ in exchange for batch $S_{j}\subseteq\{a_{1},\dots,a_{m}\}$. These updates remain entirely off-chain, enabling multiple negotiations without incurring blockchain overhead. Once the parties agree on a settlement, $\Pi.\mathsf{Lock}$ is executed: both sides upload the agreed final state to $\xi^{\mathsf{1}}$ and $\xi^{\mathsf{2}}$, where hash-locked conditions $h(\rho)$ with timeouts $T_{1}>T_{2}$ are installed to ensure atomicity. If $\mathcal{B}$ reveals the preimage $\rho$ on $\xi^{\mathsf{2}}$ before $T_{2}$, $\Pi.\mathsf{Unlock}$ delivers the agreed union of RWAs $\bigcup_{j}S_{j}$ to $\mathcal{B}$ and allows $\mathcal{S}$ to redeem the corresponding aggregated payment on $\xi^{\mathsf{1}}$ before $T_{1}$. If instead $\mathcal{B}$ does not reveal $\rho$ in time, $\Pi.\mathsf{Refund}$ is triggered: $\xi^{\mathsf{2}}$ returns the RWAs to $\mathcal{S}$ at $T_{2}$, and $\xi^{\mathsf{1}}$ returns the funds to $\mathcal{B}$ after $T_{1}$. Finally, once the session concludes, $\Pi.\mathsf{Close}$ is called to release any remaining deposits and terminate the channel.
@@
5 Security Analysis
In this section, we provide the security analysis of xRWA.
Theorem 1 (Security of xRWA).
Assume that the underlying DID and VC infrastructure is secure, meaning that no PPT adversary can cause a verifier to accept an identity or credential that was not honestly issued by an authorized issuer, except with negligible probability. Assume further that the SPV proof system is sound for transaction inclusion, and that the cross-chain channel relies on a collision-resistant hash function $h$ together with correctly enforced smart contracts operating over immutable and consistent blockchains. Under these assumptions, the xRWA protocol $\Pi_{\mathsf{xRWA}}$ achieves cross-chain authentication correctness and cross-chain settlement atomicity, and therefore satisfies RWA security as defined in Definition 6.
Proof.
Suppose there exists a PPT adversary $\mathcal{A}$ that breaks RWA security of $\Pi_{\mathsf{xRWA}}$ with non-negligible probability. By Definition 6, this means that $\mathcal{A}$ either violates cross-chain authentication correctness or violates cross-chain settlement atomicity with non-negligible probability. We treat these two cases in turn and show that each yields an adversary that contradicts one of the assumed secure components.
First, consider cross-chain authentication correctness. Assume that the event $\mathsf{AuthAcc}_{C_{t}}(a)\wedge\neg\mathsf{AuthRec}_{C_{s}}(a)$ occurs with non-negligible probability. Here $\mathsf{AuthAcc}_{C_{t}}(a)$ denotes that $C_{t}$ accepts $a$ via a cross-chain authentication claim from $C_{s}$, and $\mathsf{AuthRec}_{C_{s}}(a)$ denotes that $C_{s}$ contains a valid and confirmed authentication record for $a$. In xRWA, $C_{t}$ accepts such a claim only after verifying an SPV proof for a transaction on $C_{s}$ and checking that the embedded commitment matches a valid credential and issuer information.
We construct a simulator that uses $\mathcal{A}$ to break one of the assumed secure components. The simulator emulates honest parties and the blockchains for $\mathcal{A}$, and forwards its queries to the DID and VC system, or the SPV proof system, depending on the component being tested. Whenever $\mathcal{A}$ causes $\mathsf{AuthAcc}_{C_{t}}(a)\wedge\neg\mathsf{AuthRec}_{C_{s}}(a)$, the simulator inspects the transcript. If the SPV verifier accepts a proof for a transaction not contained in any valid block header of $C_{s}$, the simulator outputs an attack on SPV soundness. If the transaction exists but its commitment does not correspond to any honestly issued credential, or corresponds to a revoked or nonexistent credential, the simulator outputs an attack on the DID and VC infrastructure. If the credential was honest at issuance but later removed or altered while still being accepted by $C_{t}$, the simulator outputs an attack on ledger immutability. Hence, a successful violation of cross-chain authentication correctness by $\mathcal{A}$ leads to a successful attack on one of the assumed secure components with non-negligible probability, contradicting the assumptions. Therefore, the probability that $\mathcal{A}$ causes $\mathsf{AuthAcc}_{C_{t}}(a)\wedge\neg\mathsf{AuthRec}_{C_{s}}(a)$ must be negligible.
Next, consider cross-chain settlement atomicity. Assume that the event $\mathsf{Com}_{1}\oplus\mathsf{Com}_{2}$ occurs with non-negligible probability, where $\mathsf{Com}_{i}$ denotes that chain $C_{i}$ applies the settlement updates and $\oplus$ denotes exclusive or. In xRWA, all settlements are executed through the underlying cross-chain channel, which relies on the collision-resistant hash function $h$, correctly enforced smart contracts, and immutable blockchains, and which by assumption ensures atomic settlement.
We construct a simulator that uses $\mathcal{A}$ to attack the cross-chain channel. The simulator runs $\mathcal{A}$ internally, forwards its messages to the channel challenger, and uses the challenger to emulate locking, settlement, and refund behavior. If $\mathcal{A}$ produces an execution in which exactly one chain commits the settlement, the simulator outputs this as an attack on channel atomicity, contradicting the assumed security of the channel. Consequently, the probability that $\mathcal{A}$ causes $\mathsf{Com}_{1}\oplus\mathsf{Com}_{2}$ must be negligible.
Combining the two parts, any PPT adversary that breaks RWA security of $\Pi_{\mathsf{xRWA}}$ would enable an attack on one of the assumed secure components with non-negligible probability, contradicting their security. Hence, $\Pi_{\mathsf{xRWA}}$ satisfies cross-chain authentication correctness and cross-chain settlement atomicity, and therefore satisfies RWA security. ∎
@@ 6.1 Implementation
The experiments are conducted on a workstation with an Intel Core i9-13980HX CPU and 32 GB RAM, running Windows 11 (24H2). Python 3.12.10 is used to implement and benchmark the VC routines and the SPV algorithm. The reported execution times are obtained from these Python implementations under multi-threaded settings. Smart-contract experiments are performed on Remix VM (Prague) using Solidity 0.8.30 to measure gas consumption for the HTLC and cross-chain channel workflows. The entire codebase consists of more than 2,000 lines, including the Python implementations for VC/SPV and the Solidity contracts for on-chain evaluation. Our code repository is publicly available at https://github.com/ZhuHaoming2005/xRWA.
@@ 6.2 Performance Evaluation
For the cross-chain channel, handling $n$ cross-chain interactions requires only 917,253917{,}253 gas in total, since all updates are performed off-chain and only a small constant number of operations are executed on-chain. In contrast, the HTLC scheme consumes $465{,}426\times n$ gas for $n$ interactions, as each interaction must execute both the lock and unlock operations on-chain, leading to a linear increase in cost. Moreover, compared with designs that close the channel after each cross-chain settlement, our workflow allows the channel to remain open so that it can either support the next cross-chain interaction or function as an intermediate relay, similar to a payment channel network, thereby avoiding the additional costs caused by repeatedly opening and closing the channel.
In addition to the cost advantage, the channel mechanism also improves efficiency. The time overhead of cross-chain protocols is dominated by blockchain confirmation delays. Processing $n$ interactions with HTLC requires $4\times n$ on-chain operations across two blockchains, and each operation incurs confirmation latency. By contrast, the cross-chain channel performs only four on-chain operations in total, regardless of $n$. Therefore, the channel mechanism provides both lower on-chain cost and higher efficiency compared with HTLC, particularly as the number of interactions increases.